If 6th term of a GP is 2,then the product of first 11 terms of the GP is

Grade:11

1 Answers

Arun
6 years ago
T6 = a* r5 = 2
We are asked
T1*T2* T3...T11
(a)(ar)(a*r^2).....(a*r^10)
= a^11* r^55
= (a*r^5)11
= 2^11